# Gradient-Direction-Aware Density Control for 3D Gaussian Splatting

### v1

@@ Appendix C Detailed results
The detailed performance metrics across a total of 13 different scenes from the three datasets are shown in Tables 34578, and 9 while efficiency metrics are presented in Tables 6 and 10. Table 11 presents the statistical error analysis. Render quality comparisons are demonstrated in Figures 9 and 10, where significant differences are highlighted with red boxes and enlarged in the bottom-right corners. Figure 11 compares over-densification phenomena, with FPS values and the number of Gaussians annotated in the top-left corners. The comparison between nonlinear weights and linear weights is shown in the Table 12.
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.496 0.431 0.606 0.523 0.509 0.841 0.759 0.648 0.814 INGP-Base 0.491 0.450 0.649 0.574 0.518 0.855 0.798 0.818 0.890 INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 Mip-NeRF360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941 3D-GS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938 AbsGS-0004 0.782 0.613 0.870 0.784 0.626 0.920 0.908 0.929 0.944 Pixel-GS 0.766 0.627 0.862 0.784 0.638 0.929 0.921 0.936 0.951 GDAGS 0.776 0.615 0.861 0.786 0.624 0.930 0.915 0.933 0.948
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 21.912 20.097 23.495 20.661 22.248 27.594 23.624 23.420 24.669 INGP-Base 22.193 20.348 24.599 23.626 22.364 29.269 26.439 28.548 30.337 INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690 26.691 29.479 30.685 Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46 3D-GS 25.246 21.520 27.410 26.550 22.490 30.632 28.700 30.317 31.980 AbsGS-0004 25.373 21.298 27.579 26.766 22.074 31.582 28.968 31.774 32.283 Pixel-GS 25.263 21.677 27.393 26.938 22.343 31.952 29.270 31.950 32.687 GDAGS 25.309 21.408 27.465 26.960 22.034 31.820 29.015 31.951 32.277
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.506 0.521 0.386 0.503 0.540 0.4186 0.441 0.447 0.398 INGP-Base 0.487 0.481 0.312 0.450 0.489 0.301 0.342 0.254 0.227 INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176 3D-GS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205 AbsGS-0004 0.186 0.295 0.104 0.202 0.297 0.216 0.198 0.124 0.194 Pixel-GS 0.206 0.288 0.111 0.208 0.301 0.188 0.166 0.109 0.167 GDAGS 0.195 0.297 0.117 0.215 0.310 0.186 0.182 0.116 0.172
@@
Truck Train Dr Johnson Playroom Plenoxels 0.774 0.663 0.787 0.802 INGP-Base 0.779 0.666 0.839 0.754 INGP-Big 0.800 0.689 0.854 0.779 Mip-NeRF360 0.857 0.660 0.901 0.900 3D-GS 0.879 0.802 0.899 0.906 AbsGS-0004 0.886 0.820 0.900 0.907 Pixel-GS 0.886 0.826 0.887 0.905 GDAGS 0.884 0.811 0.901 0.909
@@
Truck Train Dr Johnson Playroom Plenoxels 23.221 23.221 23.142 22.980 INGP-Base 23.260 20.170 27.750 19.483 INGP-Big 23.383 20.456 28.257 21.665 Mip-NeRF360 24.912 19.523 29.140 29.657 3D-GS 25.187 21.097 28.766 30.044 AbsGS-0004 25.702 22.010 28.930 29.967 Pixel-GS 25.467 22.028 28.128 29.700 GDAGS 25.560 21.557 29.232 30.166
@@
Truck Train Dr Johnson Playroom Plenoxels 0.335 0.422 0.521 0.499 INGP-Base 0.274 0.386 0.381 0.465 INGP-Big 0.249 0.360 0.352 0.428 Mip-NeRF360 0.159 0.354 0.237 0.252 3D-GS 0.148 0.218 0.244 0.241 AbsGS-0004 0.131 0.193 0.240 0.232 Pixel-GS 0.120 0.178 0.255 0.241 GDAGS 0.143 0.208 0.238 0.233
@@
Table 11: Error analysis.
[…]
Table 12: The LPIPS comparison between nonlinear weights and linear weights. Linear weight set to $2-\mathcal{C}_{i}$.
playroom Train bicycle flowers nonlinear weights 0.233 0.208 0.195 0.297 linear weights 0.238 0.210 0.218 0.340
@@
Appendix D Limitations of GDAGS

### v2

@@ Appendix E Detailed results
The detailed performance metrics across a total of 13 different scenes from the three datasets are shown in Tables 89101213, and 14 while efficiency metrics are presented in Tables 11 and 15. Render quality comparisons are demonstrated in Figures 12 and 13, where significant differences are highlighted with red boxes and enlarged in the bottom-right corners. Figure 14 compares over-densification phenomena, with FPS values and the number of Gaussians annotated in the top-left corners.
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.496 0.431 0.606 0.523 0.509 0.841 0.759 0.648 0.814 INGP-Base 0.491 0.450 0.649 0.574 0.518 0.855 0.798 0.818 0.890 INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 Mip-NeRF360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941 3D-GS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938 AbsGS-0004 0.782 0.613 0.870 0.784 0.626 0.920 0.908 0.929 0.944 Pixel-GS 0.766 0.627 0.862 0.784 0.638 0.929 0.921 0.936 0.951 GDAGS 0.798 0.641 0.879 0.796 0.638 0.951 0.928 0.954 0.961
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 21.912 20.097 23.495 20.661 22.248 27.594 23.624 23.420 24.669 INGP-Base 22.193 20.348 24.599 23.626 22.364 29.269 26.439 28.548 30.337 INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690 26.691 29.479 30.685 Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46 3D-GS 25.246 21.520 27.410 26.550 22.490 30.632 28.700 30.317 31.980 AbsGS-0004 25.373 21.298 27.579 26.766 22.074 31.582 28.968 31.774 32.283 Pixel-GS 25.263 21.677 27.393 26.938 22.343 31.952 29.270 31.950 32.687 GDAGS 25.738 21.711 27.839 27.064 22.217 32.566 29.531 32.698 32.839
@@
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.506 0.521 0.386 0.503 0.540 0.4186 0.441 0.447 0.398 INGP-Base 0.487 0.481 0.312 0.450 0.489 0.301 0.342 0.254 0.227 INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176 3D-GS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205 AbsGS-0004 0.186 0.295 0.104 0.202 0.297 0.216 0.198 0.124 0.194 Pixel-GS 0.206 0.288 0.111 0.208 0.301 0.188 0.166 0.109 0.167 GDAGS 0.165 0.266 0.094 0.185 0.274 0.092 0.098 0.059 0.070
@@
Truck Train Dr Johnson Playroom Plenoxels 0.774 0.663 0.787 0.802 INGP-Base 0.779 0.666 0.839 0.754 INGP-Big 0.800 0.689 0.854 0.779 Mip-NeRF360 0.857 0.660 0.901 0.900 3D-GS 0.879 0.802 0.899 0.906 AbsGS-0004 0.886 0.820 0.900 0.907 Pixel-GS 0.886 0.826 0.887 0.905 GDAGS 0.886 0.819 0.901 0.909
@@
Truck Train Dr Johnson Playroom Plenoxels 23.221 23.221 23.142 22.980 INGP-Base 23.260 20.170 27.750 19.483 INGP-Big 23.383 20.456 28.257 21.665 Mip-NeRF360 24.912 19.523 29.140 29.657 3D-GS 25.187 21.097 28.766 30.044 AbsGS-0004 25.702 22.010 28.930 29.967 Pixel-GS 25.467 22.028 28.128 29.700 GDAGS 25.617 21.971 29.232 30.166
@@
Truck Train Dr Johnson Playroom Plenoxels 0.335 0.422 0.521 0.499 INGP-Base 0.274 0.386 0.381 0.465 INGP-Big 0.249 0.360 0.352 0.428 Mip-NeRF360 0.159 0.354 0.237 0.252 3D-GS 0.148 0.218 0.244 0.241 AbsGS-0004 0.131 0.193 0.240 0.232 Pixel-GS 0.120 0.178 0.255 0.241 GDAGS 0.136 0.195 0.238 0.233
@@
Appendix F Limitations of GDAGS